# Towards scalable quantum computations of atomic nuclei

### v1

@@ IV Results from quantum simulations
Figure 2: Ground state energies and state fidelities obtained using ADAPT-VQE to solve the deuteron. Results are shown as a function of the number of optimization steps $N_{\rm opt}$. We optimize the parameters $\theta_{\alpha\beta}$ for a selection of 10 operators $\hat{A}_{\alpha\beta}$ in each epoch $\alpha$, where the start of an epoch and the selection of a new set of operators is indicated by the open points. In each epoch, we use at most 100 optimization iterations to identify an optimal set of parameters. The exact ground state energy and the three-fold degenerate first excited-state energy for our Hamiltonian, computed via exact diagonalization, are indicated as dashed black and dotted gray lines, respectively.
@@ IV.1 Simulations without noise
Our exact simulation of the deuteron is summarized in Fig. 2. We show the ground-state energy, the energy difference to the exact ground state energy $E_{\mathrm{FCI}}$ as computed from an exact diagonalization [or full configuration interaction (FCI)], and the fidelity of our state $|\vec{\theta}\rangle$ as a function of the number of optimization steps.
Our initial state has an energy of 62.206 MeV, far from the exact ground state energy of $-12.874$ MeV. The corresponding fidelity with the exact ground state $\left|\psi\right\rangle$, computed as $|\left\langle\psi|\phi_{0}\right\rangle|^{2}$, is very low, specifically 0.013. In this case, our initial state is clearly too compact to quantitatively describe the weakly bound deuteron, leading to the large energy expectation value and low fidelity. Additionally, our initial state is not translationally invariant; the construction of a symmetry preserving state, as discussed in Appendix C, would require the preparation of a complicated linear combination of our initial state on all sites of the lattice. The exact ground state is of course translationally invariant, which is another contributing factor to the low fidelity of $\left|\phi_{0}\right\rangle$ with the exact ground state.
In our first two epochs, we see a substantial decrease in the ground-state energy, decreasing below the three-fold degenerate (and spurious) first excited state by the end of the second epoch. During this phase of our calculation, the fidelity increases substantially, but not quite monotonically because the optimizer spends some time exploring the local saddle points in parameter space associated with the excited states. In later epochs, as our ADAPT-VQE ansatz grows in terms of the number of exponentials $N_{\rm e}$, the energy continues to decrease, converging to within roughly 1 MeV of the exact ground state energy in 500 optimization steps with $N_{\rm e}=5$. At this point, the fidelity is also very close to 1. Further expanding and optimizing our ansatz continues to improve our state to the point where reproduction of the exact ground state and ground state energy could basically be considered exact.
@@
Figure 3 shows similar results for our exact simulation of 3He. Our initial state has an energy of 9.331 MeV, considerable closer to the exact ground state energy of $-29.468$ MeV than in the case of the deuteron. Still the initial state lies far above the first excited state and has a very small fidelity of 0.030, which we attribute to missing translational invariance and the too-compact structure of the initial state. However, the fidelity of the initial state for 3He is larger than for the deuteron, presumably because the former actually is more compact than the latter.
We find that our calculation systematically converges to the exact ground state energy and achieves essentially perfect fidelity with the exact ground state. The number of optimization steps required is larger than for the deuteron, but even after 8 epochs, we already predict an energy closer to the ground state energy than the first excited state energy. We partially attribute the longer optimization time to the larger pool of potential operators for our ADAPT-VQE ansatz due to the larger basis size ($n_{q}=3L^{3}$ rather than $2L^{3}$). However, it is important to note that we do not include any additional three-body operators in our ansatz for our 3He calculations. This is crucial, because we are able to efficiently capture the essential correlations in the three-body system using an ansatz consisting of only one- and two-body operators, suggesting that such an approach may also scale well to heavier systems. This is also intuitive, as in classical many-body methods the truncation at the normal-ordered two-body level has been demonstrated to be a very effective, controlled approximation [82, 83, 84, 85, 86].
@@ IV.2 Simulations with measurement noise
Figure 4: Differences of ground state energies of the deuteron to the exact ground state energy $E_{\mathrm{FCI}}$ of the deuteron as computed in ADAPT-VQE simulations including measurement noise for $N_{\rm shots}=1000$ (blue), 10000 (red) and without measurement noise (black). Results are shown as a function of the number of optimization steps $N_{\rm opt}$. The open points with error bars are results at the end of each optimization epoch, with uncertainties estimated as $E_{\mathrm{corr}}/\sqrt{N_{\rm shots}}$ based on the correlation energy and the number of shots. For clarity, we only show 10 % of the evaluated energies obtained during the optimization as small, filled-in points. The estimated uncertainty on the energy during the optimization is indicated by the band.
@@
We compare $N_{\rm shots}{}=1000$ (blue), 10000 (red) to the exact simulation discussed above in black. We see that in the first few epochs, the sampling uncertainty does not pose a significant challenge for the optimizer and the calculations with $N_{\rm shots}{}=1000$, 10000 both closely follow the exact calculation. Once we have converged within about 5–10 MeV of the ground state energy, the sampling uncertainty slows the optimization down slightly, but in both cases the ground state energy continues to systematically decrease. By the completion of the sixth epoch, we see that the intrinsic uncertainty from the finite number of samples limits the further optimization of our state. At this point the energy difference to the exact state is of roughly the same size as the estimated sampling uncertainty on the energy, and so additional epochs would only be able to improve the state if a larger number of samples would be used. We do not explore such adaptive optimization strategies in this work.
The results from Fig. 4 are consistent with the estimates presented in Fig. 1, as the $A=2$ system indeed requires of the order of $10^{4}$ shots to get a 1 MeV accuracy.
@@ IV.3 Practical considerations and perspectives
Figure 5: Number of $T$ (top) and sequential CNOT (bottom) gates in the circuit for our ADAPT-VQE ansatz as the number of exponentials $N_{\rm e}$ is increased for the deuteron and 3He. As the number of exponentials is increased, the ansatz grows more complicated and the circuit grows due to the added operators. At the same time, the error to the exact ground state energy ($E-E_{\mathrm{FCI}}$ plotted on the $x$ axis) is systematically decreased.
@@
Next, we consider the cost of repeated measurements on hardware. Based on the estimates in Sec. III.3 and Fig. 1 and our simulations including measurement noise, we find that reaching 1 MeV precision requires (at most) on the order of $10^{4}$$10^{6}$ measurements per optimization step. This cost scales mildly with system size and lattice size, but is already important. Based on these findings, ADAPT-VQE appears more attractive as a way to prepare an initial state for a more efficient algorithm like QPE, where the optimization would be stopped once a sufficient fidelity has been reached. We find in Figs. 2 and 3 that our ADAPT-VQE ansätze effectively produce states with high fidelity with the ground state, hinting that our scheme could serve as a basis for preparing nuclear ground states on larger lattices. For example, for the deuteron (triton) we reach a fidelity greater than 0.6 for $N_{\rm e}=3$ ($N_{\rm e}=6$). We refer the reader to Ref. [54], where the cost of simulating the time evolution of a similar Hamiltonian has been estimated, both using Trotter-Suzuki product formula and qubitization.
@@ V Summary
Using ADAPT-VQE, we performed simulations of quantum computations of the deuteron and 3He with a simple Hamiltonian from pionless effective field theory including two- and three-nucleon forces. We found that our calculations are able to systematically converge towards the exact ground state results as we expand the complexity of the variational ansatz used in ADAPT-VQE. Our calculations also performed well in simulations involving a finite number of stochastic measurements of the quantum circuit.
@@ Acknowledgements.
This work was supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Nos. DE-FG02-96ER40963 and DE-SC0021642, by the Quantum Science Center, a National Quantum Information Science Research Center of the U.S. Department of Energy, and by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the U.S. Department of Energy. Oak Ridge National Laboratory is supported by the Office of Science of the Department of Energy under contract No. DE-AC05-00OR22725. This research used resources of the Oak Ridge Leadership Computing Facility located at Oak Ridge National Laboratory. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). OK is supported by CERN through the CERN Quantum Technology Initiative.
@@ Appendix A Hamiltonian
Table 1: Ground-state energies of 2H and 3He (in MeV) from exact diagonalizations on lattices of extent $L$, lattice spacing $a=2$ fm, two-body coupling $v=-9.0$, and three-body coupling $w=10.8$.
@@ Appendix B ADAPT-VQE operator pool
For the pool of operators for our ADAPT-VQE calculations, we focus on a pool of operators producing orthogonal transformations. We found that for our real-valued Hamiltonian this is more effective at generating appropriate transformations that lower the energy expectation value than general unitary transformations. This can be understood through the following argument.
@@
onto the state $|j\rangle$. Here, $\theta$ and $\phi$ are both real numbers. The operator (17) generates an orthogonal transformation (a Givens rotation), while the operator (18) generates a unitary transformation via the exponentiation of a Hermitian generator multiplied with a purely imaginary number. In the first case we have
@@
while in the second
@@
For real-symmetric Hamiltonians $\langle j|\hat{H}|k\rangle=\langle k|\hat{H}|j\rangle$ and the last line in Eq. (B) vanishes. Thus the energy will fulfill $E(\phi)<E(0)$ only if $\langle k|\hat{H}|k\rangle<\langle j|\hat{H}|j\rangle$. In contrast, the operator $\hat{A}(\theta)$ lowers the energy when acting on the state $|j\rangle$ for any $\langle k|\hat{H}|j\rangle\neq 0$. To see this, one can take $|\theta|\ll 1$ in Eq. (B) and choose its sign such that $\theta\langle k|\hat{H}|j\rangle<0$. We clearly see that in this case it is generally easier to reduce the energy of our initial state through the application of an appropriately parametrized orthogonal transformation.
@@
[…]
@@ Appendix C Translationally invariant initial state
In the last step, we employed the angles (28). Clearly, the state $|\psi_{0}\rangle$ is a coherent superposition of states where a deuteron is on each lattice site. It is also a correlated state. The effort of its preparation is ${\cal O}(n)$. In practice, however, the preparation can only be implemented approximately, because Eq. (27) requires a Trotterization. As is seems questionable to implement the exact symmetry corresponding to translational invariance only approximately, we did not pursue this further.

### v2

@@ IV Results from quantum simulations
Figure 2: Ground state energies and state fidelities obtained using ADAPT-VQE to solve the deuteron. Results are shown as a function of the number of optimization steps $N_{\rm opt}$. We optimize the parameters $\theta_{\alpha\beta}$ for a selection of 10 operators $\hat{A}_{\alpha\beta}$ in each epoch $\alpha$, where the start of an epoch and the selection of a new set of operators is indicated by the open points. In each epoch, we use at most 100 optimization iterations to identify an optimal set of parameters. The exact ground state energy and the threefold degenerate first excited-state energy for our Hamiltonian, computed via exact diagonalization, are indicated as dashed black and dotted gray lines, respectively.
@@ IV.1 Simulations without noise
Our exact simulation of the deuteron is summarized in Fig. 2. We show the ground-state energy, the energy difference with the exact ground state energy $E_{\mathrm{FCI}}$ as computed from an exact diagonalization [or full configuration interaction (FCI)], and the fidelity of our state $|\vec{\theta}\rangle$ as a function of the number of optimization steps.
Our initial state has an energy of 62.206 MeV, far from the exact ground state energy of 12.874-12.874 MeV. The corresponding fidelity with the exact ground state $\left|\psi\right\rangle$, computed as $|\left\langle\psi|\phi_{0}\right\rangle|^{2}$, is very low, specifically 0.013. In this case, our initial state is clearly too compact to quantitatively describe the weakly bound deuteron, leading to the large energy expectation value and low fidelity. Additionally, our initial state is not translationally invariant; the construction of a symmetry preserving state, as discussed in Appendix C, would require the preparation of a complicated linear combination of our initial state on all sites of the lattice. The exact ground state is of course translationally invariant, which is another contributing factor to the low fidelity of $\left|\phi_{0}\right\rangle$ with the exact ground state.
In our first two epochs, we see a substantial decrease in the ground-state energy, decreasing below the threefold degenerate (and spurious) first excited state by the end of the second epoch. During this phase of our calculation, the fidelity increases substantially, but not quite monotonically because the optimizer spends some time exploring the local saddle points in parameter space associated with the excited states. In later epochs, as our ADAPT-VQE ansatz grows in terms of the number of exponentials $N_{\rm e}$, the energy continues to decrease, converging to within roughly 1 MeV of the exact ground state energy in 500 optimization steps with $N_{\rm e}=5$. At this point, the fidelity is also very close to 1. Further expanding and optimizing our ansatz continues to improve our state to the point where reproduction of the exact ground state and ground state energy could basically be considered exact.
@@
Figure 3 shows similar results for our exact simulation of 3He. Our initial state has an energy of 9.331 MeV, considerable closer to the exact ground state energy of 29.468-29.468 MeV than in the case of the deuteron. Still the initial state lies far above the first excited state and has a very small fidelity of 0.030, which we attribute to missing translational invariance and the too-compact structure of the initial state. However, the fidelity of the initial state for 3He is larger than for the deuteron, presumably because the former actually is more compact than the latter.
We find that our calculation systematically converges to the exact ground state energy and achieves essentially perfect fidelity with the exact ground state. The number of optimization steps required is larger than for the deuteron, but even after eight epochs, we already predict an energy closer to the ground state energy than the first excited state energy. We partially attribute the longer optimization time to the larger pool of potential operators for our ADAPT-VQE ansatz due to the larger basis size ($n_{q}=3L^{3}$ rather than $2L^{3}$). However, it is important to note that we do not include any additional three-body operators in our ansatz for our 3He calculations. This is crucial, because we are able to efficiently capture the essential correlations in the three-body system using an ansatz consisting of only one- and two-body operators, suggesting that such an approach may also scale well to heavier systems. This is also intuitive, as in classical many-body methods the truncation at the normal-ordered two-body level has been demonstrated to be a very effective, controlled approximation [102, 103, 104, 105].
We can contrast this to the work of Pérez-Obiol et al. [10], where they performed quantum shell-model computations of 20,22,24Ne of four, six, and eight interacting nucleons, respectively, using 24 qubits. Their ADAPT-VQE simulations reached precisions of 2% after including 167–345 layers in the ADAPT-VQE ansatz. Our calculation of 3He reaches the same precision at 30 layers, clearly demonstrating that a simpler ansatz with a significantly smaller circuit depth is sufficient in our case. This is due to the simplicity of the Hamiltonian and the natural expression of physically relevant operators (e.g., nearest-neighbor hopping operators) on the position-space lattice. A more expressive operator pool could potentially even further simplify the required ansatz and reduce the circuit depth, as found for shell-model calculations in Ref. [15].
@@ IV.2 Simulations with measurement noise
Figure 4: Differences of ground state energies of the deuteron with the exact ground state energy $E_{\mathrm{FCI}}$ of the deuteron as computed in ADAPT-VQE simulations including measurement noise for $N_{\rm shots}=1000$ (blue), 10000 (red) and without measurement noise (black). Results are shown as a function of the number of optimization steps $N_{\rm opt}$. The open points with error bars are results at the end of each optimization epoch, with uncertainties estimated as $E_{\mathrm{corr}}/\sqrt{N_{\rm shots}}$ based on the correlation energy and the number of shots. For clarity, we only show 10% of the evaluated energies obtained during the optimization as small, filled-in points. The estimated uncertainty on the energy during the optimization is indicated by the band.
@@
We compare $N_{\rm shots}{}=1000$ (blue), 10000 (red) with the exact simulation discussed above in black. We see that in the first few epochs, the sampling uncertainty does not pose a significant challenge for the optimizer and the calculations with $N_{\rm shots}{}=1000$, 10000 both closely follow the exact calculation. Once we have converged within about 5–10 MeV of the ground state energy, the sampling uncertainty slows the optimization down slightly, but in both cases the ground state energy continues to systematically decrease. By the completion of the sixth epoch, we see that the intrinsic uncertainty from the finite number of samples limits the further optimization of our state. At this point the energy difference to the exact state is of roughly the same size as the estimated sampling uncertainty on the energy, and so additional epochs would only be able to improve the state if a larger number of samples would be used. We do not explore such adaptive optimization strategies in this work.
The results from Fig. 4 are consistent with the estimates presented in Fig. 1 because the $A=2$ system indeed requires on the order of $10^{4}$ shots to get a 1 MeV accuracy.
@@ IV.3 Practical considerations and perspectives
Figure 5: Number of $T$ (top) and sequential CNOT (bottom) gates in the circuit for our ADAPT-VQE ansatz as the number of exponentials $N_{\rm e}$ is increased for the deuteron and 3He. As the number of exponentials is increased, the ansatz grows more complicated and the circuit grows due to the added operators. At the same time, the error to the exact ground state energy ($E-E_{\mathrm{FCI}}$ plotted on the $x$ axis) is systematically decreased. We only show points with $N_{\rm e}\geq 1$, where the number of $T$ and CNOT gates is larger than zero.
@@
Next, we consider the cost of repeated measurements on hardware. Based on the estimates in Sec. III.3 and Fig. 1 and our simulations including measurement noise, we find that reaching 1 MeV precision requires (at most) on the order of $10^{4}$$10^{6}$ measurements per optimization step. This cost scales mildly with system size and lattice size, but is already important. Based on these findings, ADAPT-VQE appears more attractive as a tool for preparing approximate initial states, since the optimization need not reach unit fidelity with the ground state. Instead, once a sufficient overlap (e.g., 10%) is obtained, the state can be passed to fault-tolerant quantum algorithms with guarantees, such as QPE or other filtering methods, which can then extract the ground-state component. We find in Figs. 2 and 3 that our ADAPT-VQE ansätze effectively produce states with high fidelity with the ground state, hinting that our scheme could serve as a basis for preparing nuclear ground states on larger lattices. For example, for the deuteron (triton) we reach a fidelity greater than 0.6 for $N_{\rm e}=3$ ($N_{\rm e}=6$). We refer the reader to Ref. [58], further improved in Ref. [107], where the cost of simulating the time evolution of a similar Hamiltonian has been estimated, both using Trotter-Suzuki product formula and qubitization.
@@ V Summary
Using ADAPT-VQE, we performed simulations of quantum computations of the deuteron and 3He with a simple Hamiltonian from pionless effective field theory including two- and three-nucleon forces. We found that our calculations are able to systematically converge toward the exact ground state results as we expand the complexity of the variational ansatz used in ADAPT-VQE. Our calculations also performed well in simulations involving a finite number of stochastic measurements of the quantum circuit.
@@
Data availability
The data shown in the figures and table of this work are available in Ref. [108].
@@ Acknowledgements.
This work was supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award No. DE-FG02-96ER40963 and No. DE-SC0021642, by the Quantum Science Center, a National Quantum Information Science Research Center of the U.S. Department of Energy, and by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the U.S. Department of Energy. Oak Ridge National Laboratory is supported by the Office of Science of the Department of Energy under contract No. DE-AC05-00OR22725. This research used resources of the Oak Ridge Leadership Computing Facility located at Oak Ridge National Laboratory. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). OK is supported by CERN through the CERN Quantum Technology Initiative. This manuscript has been authored in part by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
@@ Appendix A Hamiltonian
Table 1: Ground-state energies of 2H and 3He (in MeV) from exact diagonalizations on lattices of extent $L$, lattice spacing $a=2$ fm, two-body coupling v=9.0v=-9.0, and three-body coupling $w=10.8$.
@@ Appendix B ADAPT-VQE operator pool
For the pool of operators for our ADAPT-VQE calculations, we focus on a pool of operators producing orthogonal transformations. We found that for our real-valued Hamiltonian this is more effective at generating appropriate transformations that lower the energy expectation value than general unitary transformations. This can be understood through the following argument:
@@
onto the state $|j\rangle$. Here, $\theta$ and $\phi$ are both real numbers. The operator (17) generates an orthogonal transformation (a Givens rotation), while the operator (18) generates a unitary transformation via the exponentiation of a Hermitian generator multiplied with a purely imaginary number. In the first case, we have
@@
while in the second case, we have
@@
For real-symmetric Hamiltonians $\langle j|\hat{H}|k\rangle=\langle k|\hat{H}|j\rangle$ and the last line in Eq. (B) vanishes. Thus the energy will satisfy $E(\phi)<E(0)$ only if $\langle k|\hat{H}|k\rangle<\langle j|\hat{H}|j\rangle$. In contrast, the operator $\hat{A}(\theta)$ lowers the energy when acting on the state $|j\rangle$ for any $\langle k|\hat{H}|j\rangle\neq 0$. To see this, one can take $|\theta|\ll 1$ in Eq. (B) and choose its sign such that $\theta\langle k|\hat{H}|j\rangle<0$. We clearly see that in this case it is generally easier to reduce the energy of our initial state through the application of an appropriately parametrized orthogonal transformation.
@@
$\hat{A}(\theta)=\exp\left(\theta\,\hat{a}^{\dagger}_{\mathbf{l}\tau_{1}s_{1}}\hat{a}^{\dagger}_{\mathbf{k}\tau_{2}s_{2}}\hat{a}_{\mathbf{k}^{\prime}\tau_{2}s_{2}}\hat{a}_{\mathbf{l}^{\prime}\tau_{1}s_{1}}-\text{H.c.}\right)\ .$ (24)
@@ Appendix C Translationally invariant initial state
In the last step, we employed the angles (28). Clearly, the state $|\psi_{0}\rangle$ is a coherent superposition of states where a deuteron is on each lattice site. It is also a correlated state. The effort of its preparation is O(n){O}(n). In practice, however, the preparation can only be implemented approximately, because Eq. (27) requires a Trotterization. As is seems questionable to implement the exact symmetry corresponding to translational invariance only approximately, we did not pursue this further.